Grade 12Mechanics

a 5.0 kg mass hamgs at the end of a cord. find the tension in the cord if the acceleration of the mass is a.) 1.5 meter per second square up, b.) 1.5 meter per second down, c.) 9.8 meter per second down.

madona banhan
8 Years agoGrade 12

1 Answer

Arun
8 Years ago
If the mass is hanging from only the cord then it would seem that its acceleration must be either upward or downward. The force to accelerate the mass = 5 x 1.5 = 7.5 N 
The tension in the cord is then the weight of the mass + or - the acceleration force 
= 49.05 + 7.5 = 56.55 N if acceleration is upward and 
= 49.05 - 7.5 = 41.55 N if acceleration is downward
Now if acceleration is 9.8 m/sec² downward
Tension = 0
Regards
Arun (askIITians forum expert)